Consider the solubilities of a particular solute at two different temperatures.

Temperature (∘C) Solubility (g/100 g H2O)
20.0 25.4
30.0 68.9

Suppose a saturated solution of this solute was made using 66.0 g H2O at 20.0 °C. How much more solute can be added if the temperature is increased to 30.0 ∘C?

mass: g
Answer #1

We know ,

At 20oC,

solubity is 25.4g of solute per 100g H2O which means

100g H2O has 25.4g of solute

so, 66g H2O has 66x25.4/100 = 16.764 g of solute

Now at 30oC,

solubility is 68.9g of solute per 100g   H2O

which implies 100g  H2O can have 68.9g of solute

Thus, 66g of H2O can have  66x68.9/100 = 45.474 g of solute

so, amount of solute which can be further added to the solution after raising the temperature to 30oC

= 45.474-16.764 = 28.71 g